OpenStudy (haleyelizabeth2017):

f(x)=3x+12 a. determine the inverse of this function and name it g(x). b. use composite functions to show that these functions are inverses. c. evaluate f(g(-2)). Explain: what is the domain?

OpenStudy (haleyelizabeth2017):

(I am putting y instead of f(x) so it is easier...) is the inverse x=3y+12 -3y=-x+12 y=1/3x-4 correct? I'm not sure what to do for b....

OpenStudy (anonymous):

looks good to me

OpenStudy (haleyelizabeth2017):

okay

OpenStudy (anonymous):

b. use composite functions to show that these functions are inverses. put \(f(x)=3x+12,g(x)=\frac{1}{3}x-4\) then compute \[g(f(x))\] and see that you get \(x\) as the final answer

OpenStudy (haleyelizabeth2017):

so do I plug it in for x? I'm confused :(

OpenStudy (anonymous):

do you know how to write a composite function?

OpenStudy (haleyelizabeth2017):

sorta........but right now my brain is fried from too much math and homework that it isn't wanting to work....so g(x)=1/3(3x+12)+4 correct?

OpenStudy (anonymous):

\[g(f(x))=g(3x+12)=\frac{1}{3}(3x+12)-4\] yes, except you made a typo and had a \(+4\) at the end, it should be \(-4\)

OpenStudy (haleyelizabeth2017):

then g(x)=x+8 ?

OpenStudy (anonymous):

exactly because you made that mistake

OpenStudy (haleyelizabeth2017):

whoopsies!

OpenStudy (anonymous):

now you get just \(x\) right?

OpenStudy (haleyelizabeth2017):

actually it is just g(x)=x then

OpenStudy (haleyelizabeth2017):

yep

OpenStudy (anonymous):

no dear, it is not \(g(x)=x\) it is \(g(f(x))=x\)

OpenStudy (anonymous):

but that is probably what you meant

OpenStudy (haleyelizabeth2017):

oh! okay

OpenStudy (haleyelizabeth2017):

sorry.....my mistake :)

OpenStudy (haleyelizabeth2017):

okay so what do I do next?

OpenStudy (anonymous):

go have a beer? oh no, still need c. evaluate f(g(-2)). Explain: what is the domain

OpenStudy (haleyelizabeth2017):

lol

OpenStudy (anonymous):

\[f(g(-2))=-2\] because the functions are invereses

OpenStudy (haleyelizabeth2017):

okay

OpenStudy (anonymous):

i do not know what "explain what is the domain" means it is not english

OpenStudy (haleyelizabeth2017):

wait........I'm confused on how that is possible....sorry! lol

OpenStudy (anonymous):

nor is it math

OpenStudy (anonymous):

how is what possible?

OpenStudy (haleyelizabeth2017):

oh never mind lol

OpenStudy (haleyelizabeth2017):

well I just need to find the domain and explain how I get that....I think

OpenStudy (anonymous):

you just proved that \[g(f(x))=x\] for any \(x\) it has to be true for \(-2\) that \[g(f(-2))=-2\]

OpenStudy (anonymous):

the domain of a polynomial function is all real numbers unless you specify otherwise in this case you have a line, so the range is all real numbers as well

OpenStudy (haleyelizabeth2017):

okay thank you so much!!!!

OpenStudy (anonymous):

yw